Ten ounces of fruit drink is 90% juice. How many ounces of water must be added to dilute the micture to 50% fruite juice?
Answer by checkley77(12844)

.90*10=.50(10+X)
9=5+.5X
.5X=9-5
.5X=4
X=4/.5
X=8 OUNCES OF WATER IS USED.
PROOF:
.90*10=.5(10+8)
9=.5*18
9=9